OpenStudy (anonymous):

Solve the absolute value equation. Graph the solution. 2|3x-5|-8 =8

jigglypuff314 (jigglypuff314):

first isolate the absolute values so add 8 to both sides so that 2|3x - 5| = 16 then divide both sides by 2 /2 /2 so |3x - 5| = 8 then understand that |a| = b can be separated into a = b and a = -b so |3x - 5| = 8 can be separated into 3x - 5 = 8 and 3x - 5 = -8 you would then solve for x for each :)